OpenStudy (anonymous):

Michael has a collection of action figures.Of his action figures 3/8 of them are Marvel characters. Of the Marvel action figures,2/5 of them were made in Taiwan.What fraction of Michael's action figures are Marvel characters and made in Taiwan??

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

multiply

OpenStudy (anonymous):

Really?Why?

OpenStudy (misty1212):

i am sure you can figure out why if \(\frac{3}{8}\) of my boyfriends are under the age of 21, and \(\frac{2}{5}\) of those those are cute, then only \(\frac{3}{8}\times \frac{2}{5}\) are both under 21 AND cute

OpenStudy (anonymous):

Oh!!Ok thanx!!